OpenStudy (anonymous):

The measures of the angles of a triangle are in the ratio 2:4:6. What type of triangle is it? a. It cannot be a triangle b. right c. obtuse d. acute

OpenStudy (mathstudent55):

2x + 4x + 6x = 180 What is x?

OpenStudy (anonymous):

2x+4x+6x=180 12x=180 x=15

OpenStudy (mathstudent55):

Great. The angles are 2x = 2(15) = ? 4x = 4(15) = ? 6x = 6(15) = ? What are the three angle measures?

OpenStudy (anonymous):

All right so... 2(15)=30 4(15)=60 6(15)=90

OpenStudy (anonymous):

It's a 30 60 90 triangle!

OpenStudy (mathstudent55):

Ok. Now read the choices and answer the question.

OpenStudy (anonymous):

It would be b) right Am I right? :)

OpenStudy (mathstudent55):

You certainly are. Great job.

OpenStudy (anonymous):

Thanks!

OpenStudy (mathstudent55):

yw